# Analytics, have some humility: a statistical view of fourth-down decision making

### v4

@@ Abstract
The standard mathematical approach to fourth-down decision making in American football is to make the decision that maximizes estimated win probability. Win probability estimates arise from machine learning models fit from historical data. These models, however, are overfit high-variance estimators, exacerbated by the highly correlated nature of football play-by-play data. Thus, it is imperative to knit uncertainty quantification into the fourth-down decision procedure; we do so using bootstrapping. We find that uncertainty in the estimated optimal fourth-down decision is far greater than that currently expressed by sports analysts in popular sports media. Our contribution is a major advance in fourth-down strategic decision making: far fewer fourth-down decisions are as obvious as analysts claim.
@@ 1 Introduction
The two most widely used value functions by analysts of American football are win probability and expected points (Yurko et al.,, 2018). Win probability (${\mathsf{WP}}$) measures the probability that the team with possession at the current game-state wins the game. Expected points (${\mathsf{EP}}$) measures the expected value of the net number of points of the next score in the game, relative to the team with possession, given the current game-state. The most prominent example of analysts using these value functions to dictate in-game strategy is fourth-down decision making. On fourth down, a football coach has three choices: go for it (${\mathsf{Go}}$), attempt a field goal (${\mathsf{FG}}$), or punt the ball (${\mathsf{Punt}}$). Previous work by Romer, (2006) and Burke, 2009b suggest making the decision that maximizes estimated ${\mathsf{EP}}$. Yet a team’s goal is to win the game, not score more points on average, so ${\mathsf{EP}}$ is the wrong objective function. Hence, modern approaches by Baldwin, (2021) and Burke, 2009a suggest making the decision that maximizes estimated ${\mathsf{WP}}$. These analyses found that National Football League (NFL) coaches are too conservative on fourth down; they often settle for kicks even when they should go for it.
@@
The ${\mathsf{WP}}$ estimates used for fourth-down decision making typically arise from statistical models fit from historical data. Given the results of the entire recent play-by-play history of football, these models fit the relationship between certain game-state variables using data-driven regression or machine learning approaches. However, these models are overfit high-variance estimators, exacerbated by the highly correlated nature of football play-by-play data. Because every game has only one winner, the binary win/loss response values are not independent, as all plays from the same game share the exact same draw of the win/loss outcome. This reduces the effective sample size and inflates standard errors.
Traditional decision procedures, based solely on effect size, ignore the uncertainty inherent in estimating ${\mathsf{WP}}$ from noisy and highly correlated historical data. This uncertainty should percolate into the fourth-down decision procedure, but popular approaches currently treat win probabilities as known quantities rather than estimates. Trusting the resulting ${\mathsf{WP}}$ point estimates to dictate fourth-down decision making without accounting for uncertainty in these estimates leads to overconfident and potentially sub-optimal decisions.
@@
Our focus is not to “fix” win probability models by adjusting for additional covariates or reducing model bias. Rather, we shed light on the high variance nature of estimating win probability and show that such estimates are rife with uncertainty. We use bootstrapping to quantify uncertainty in fourth-down recommendations and recommend a fourth-down decision when we are confident it has higher ${\mathsf{WP}}$ than all other decisions. Our contribution to the literature is a major advance in fourth-down strategic decision making: far fewer fourth-down decisions are as obvious as analysts claim. A substantial proportion of fourth-down decisions feature ${\mathsf{WP}}$ recommendations subject to large uncertainty. Thus, we ask football analysts to have some humility: for many game-states, there is simply not enough data to trust ${\mathsf{WP}}$ point estimates.
@@
The remainder of this paper is organized as follows. In Section 2 we detail the traditional fourth-down decision procedure. Then in Section 3 we incorporate uncertainty quantification into the fourth-down decision procedure. In Section 4 we present our main findings and we conclude in Section 5.
@@ 2 The traditional fourth-down decision procedure
The standard mathematical approach to fourth-down decision making is to make the decision that maximizes estimated win probability. In this Section we explore the traditional decision procedure and detail how analysts commonly estimate win probability today. We begin with a brief overview of our historical dataset of football plays in Section 2.1. Then in Section 2.2 we illustrate the traditional decision process for example plays. In Section 2.3 we describe how analysts estimate a team’s fourth-down win probability if it punts, kicks a field goal, or attempts a conversion. Fourth-down win probability is typically estimated in terms of first-down win probability, so in Section 2.4 we describe well known open source first-down win probability models.
@@ 2.1 Data
We access every NFL play from 1999 to 2022 using the ${\mathsf{R}}$ package nflFastR (Carl and Baldwin,, 2022). Each play includes variables that are relevant to estimating win probability that describe the context of the play, such as the score differential, time remaining, yardline, down, yards to go, etc. (see Table 3 in Appendix A for descriptions of relevant variables). Note that yardline is a real number in $\{0,1,...,99,100\}$ measuring the number of yards from the opponent’s endzone, where 00 represents a touchdown and $100$ indicates a safety.
In order to avoid data bleed in constructing the various decision transition models (Appendix C) relevant for fourth-down decision making, we hold out plays from 1999 to 2005. Additionally, we hold out plays from 2022 to illustrate our decision procedure on out-of-sample example plays in Section 4.1. We are then left with a primary dataset of $600,825$ plays, with $240,380$ first-down plays and $4,101$ non-tied games from 2006 to 2021, henceforth referred to as the “observed play-by-play football dataset.” We fit our win probability models using this dataset. The code for this study, which includes code to scrape the dataset used in this study, is publicly available on Github.111 https://github.com/snoopryan123/fourth_down
@@ 2.2 Example plays
The traditional fourth down decision procedure is based solely on effect size: make the decision that maximizes estimated win probability. Further, the strength of a decision is proportional to the estimated gain in win probability by making that decision. Notably, Baldwin, (2021) and Burke, 2009a employ this decision procedure and post their fourth-down recommendations on Twitter.222 Baldwin’s posts his fourth-down recommendations at @ben_bot_baldwin and Burke posts his at @bburkeESPN. We illustrate this procedure through two example plays.
Example play 1. First, we consider an example of Baldwin’s fourth-down decision charts, which is based on a model described in Section 2.4. Figure 1 illustrates Baldwin’s decision-making charts for a fourth-down play in which the Colts had the ball against the Patriots in Week 10 of 2023. Baldwin views ${\mathsf{Go}}$ as a “strong” decision because he estimates that going for it provides a $3.6\%$ gain in win probability over attempting a field goal.
Figure 1: Baldwin’s decision charts for example play 1. Baldwin views ${\mathsf{Go}}$ as a “strong” decision because he estimates that going for it provides a $3.6\%$ gain in win probability over attempting a field goal.
Example play 2. Next, we give an example of Burke’s decision chart, which is based on a proprietary model. Figure 2 illustrates Burke’s fourth-down decision chart for a play from the 2023 NFC Championship game. The chart visualizes the estimated optimal decision (color) according to effect size as a function of yardline ($x$-axis) and yards to go ($y$-axis), holding the other game-state variables constant. Burke views ${\mathsf{Go}}$ as the right decision because the yellow dot (representing the actual play’s yardline and yards to go) lies squarely in the red region and is far from the decision boundary.
Figure 2: Burke’s decision boundary chart for example play 2. The chart visualizes the estimated optimal decision according to effect size (color) as a function of yardline ($x$-axis) and yards to go ($y$-axis), holding the other game-state variables constant. Burke views ${\mathsf{Go}}$ as the right decision because the yellow dot (representing the actual play’s yardline and yards to go) lies squarely in the red region and is far from the decision boundary.
@@ 2.3 Estimating the win probability of a fourth-down decision
An offense has three possible decisions to make on fourth down: it can punt the ball, kick a field goal, or attempt a conversion (colloquially known as “going for it”), denoted ${\mathsf{Punt}}$, ${\mathsf{FG}}$, and ${\mathsf{Go}}$, respectively. To employ the traditional fourth-down decision procedure, we need to estimate a team’s fourth-down win probability for each decision as a function of the game-state.
In Appendix B we estimate the win probability of a fourth-down decision in $\{{\mathsf{Punt}},{\mathsf{FG}},{\mathsf{Go}}\}$ in terms of first-down win probability and decision transition probabilities (punt outcome distribution, field goal success probability, and conversion probability). We can do this because a first down always follows a fourth down. For instance, success on a conversion attempt results in a first down for the team with possession and failure results in a first down for the other team. Thus, ${\mathsf{Go}}$ win probability on fourth down is a function of conversion success probability and first-down win probability. Similarly, ${\mathsf{FG}}$ win probability is a function of field goal success probability and first-down win probability. ${\mathsf{Punt}}$ win probability is a function of punt outcome and and first-down win probability. We estimate these decision transition probabilities in Appendix C.
We detail first-down win probability estimation in the main text, relegating the remainder of fourth-down win probability modeling to the Appendix, because it is the most difficult component of fourth-down decision making. Whereas decision transition models are fit from i.i.d. play-level outcomes and so have a relatively large effective sample size, first-down win probability models are fit from noisy, clustered, and highly correlated game-level outcomes. There are just $4,101$ games in our dataset, providing a relatively small effective sample size for statistical first-down win probability models. On this view, first-down win probability forms the crux of the fourth-down decision framework.
2.4 Estimating the win probability of a first down
@@
Win probability estimates arise broadly from one of two classes of models, probabilistic state-space models or statistical models. On one hand, state-space models simplify the game of football into a series of transitions between game-states. Transition probabilities are estimated from play-level data and are then propagated into win probability by simulating games. When implemented correctly, these models are sensible ways to estimate ${\mathsf{WP}}$. However, they are difficult in practice, as they require: a careful encoding of the convoluted rules of football into a set of states and the actions between those states, careful estimation of transition probabilities, and enough computing power to run enough simulated games to achieve desired granularity. Each of these are nontrivial.
@@
On the other hand, statistical models are fit entirely from historical data. Given the results of a set of observed football plays, statistical models fit the relationship between certain game-state variables using data-driven regression or machine learning approaches. These models are widely used today in the football analytics community because of publicly available play-by-play data (e.g., nflFastR (Carl and Baldwin,, 2022)) and accessible off-the-shelf machine learning models (e.g., ${\mathsf{XGBoost}}$ (Chen and Guestrin,, 2016)). Additionally, due to a perceived abundance of data, flexible machine learning models are viewed as more “trustworthy” than previous mathematical models that make more simplifying assumptions. For these reasons, the open source win probability models used today for fourth-down decision making are statistical / machine learning models. Hence, in this paper we focus on statistical / machine learning win probability models.
@@
Each of these win probability models are statistical models fit from an historical play-by-play dataset. This dataset includes plays from all downs; these models fit first-down win probability by including down as a covariate. We find that a model fit from just first-down plays better predicts first-down win/loss (see Appendix D), which is what we care about for fourth-down decision making. For the decision procedure used in the remainder of this paper, we use a first-down ${\mathsf{XGBoost}}$ win probability model fit from just first downs as described in Appendix D.

### v5

@@ Abstract
The standard mathematical approach to fourth-down decision making in American football is to make the decision that maximizes estimated win probability. Win probability estimates arise from machine learning models fit from historical data. These models attempt to capture a nuanced relationship between a noisy binary outcome variable and game-state variables replete with interactions and non-linearities from a finite dataset of just a few thousand games. Thus, it is imperative to knit uncertainty quantification into the fourth-down decision procedure; we do so using bootstrapping. We find that uncertainty in the estimated optimal fourth-down decision is far greater than that currently expressed by sports analysts in popular sports media.
@@ 1 Introduction
The two most widely used value functions by analysts of American football are win probability and expected points (Yurko et al.,, 2018). Win probability (${\mathsf{WP}}$) measures the probability that the team with possession at the current game-state wins the game. Expected points (${\mathsf{EP}}$) measures the expected value of the net number of points of the next score in the game, relative to the team with possession, given the current game-state. The most prominent example of analysts using these value functions to dictate in-game strategy is fourth-down decision making. On fourth down, a football coach has three choices: go for it (${\mathsf{Go}}$), attempt a field goal (${\mathsf{FG}}$), or punt the ball (${\mathsf{Punt}}$). Previous work by Romer, (2006) and Burke, (2009) suggest making the decision that maximizes estimated ${\mathsf{EP}}$. Yet a team’s goal is to win the game, not score more points on average, so ${\mathsf{EP}}$ is the wrong objective function. Hence, modern approaches by Baldwin, (2021) and Burke111 Burke releases fourth-down decision recommendations with the ESPN analytics team, and their recommendation algorithm is proprietary. suggest making the decision that maximizes estimated ${\mathsf{WP}}$. Each of analyses found that National Football League (NFL) coaches are too conservative on fourth down; they often settle for kicks even when they should go for it.
@@
The win probability estimates used for fourth-down decision making typically arise from statistical models fit from historical data. Given the play-by-play results of the entire recent history of football, these models fit the relationship between a binary win/loss outcome variable and certain game-state variables using data-driven regression or machine learning approaches. Analysts then recommend the optimal decision according to the model. This approach overlooks the uncertainty inherent in estimating win probability with limited data. The win/loss outcome variable is noisy and there are only a few thousand games in the dataset that produce an outcome. The outcome variable, moreover, is not independent across plays because every non-tied game has only one winner, which reduces the effective sample size of the dataset. Therefore, it is critical to quantify uncertainty in win probability estimates. This uncertainty should percolate into the fourth-down decision procedure.
@@
Our focus is not to “fix” win probability models by adjusting for additional covariates or reducing model bias. Rather, we shed light on the high variance nature of estimating win probability and show that such estimates are subject to nontrivial uncertainty. We use bootstrapping to quantify uncertainty in fourth-down recommendations and recommend a decision when we are confident it has higher ${\mathsf{WP}}$ than all other decisions. We find that far fewer fourth-down decisions are as obvious as previously thought. Thus, we ask football analysts to have some humility: for many game-states, there is simply not enough data to trust ${\mathsf{WP}}$ point estimates.
@@
The remainder of this paper is organized as follows. In Section 2 we estimate win probability and detail the traditional fourth-down decision procedure. In Section 3 we knit uncertainty quantification into the decision procedure. In Section 4 we present our main findings and we conclude in Section 5.
@@ 2 The traditional fourth-down decision procedure
The standard fourth-down decision procedure involves making the decision that maximizes estimated win probability. In this section we overview this procedure and detail how analysts commonly estimate win probability today. We begin with a brief overview of our historical dataset of football plays in Section 2.1. Then in Section 2.2 we illustrate the traditional decision process through example plays. In Section 2.3 we estimate the win probability of a fourth-down decision. These estimates are functions of first-down win probability and decision transition probabilities, which we estimate in Sections 2.4 and 2.5, respectively.
@@ 2.1 Data
We access every NFL play from 1999 to 2022 using the ${\mathsf{R}}$ package nflFastR (Carl and Baldwin,, 2022). Each play includes variables that describe the context of the play, which are relevant to estimating win probability, such as the score differential, time remaining, yardline, down, yards to go, etc. (see Table 2 in Appendix A for descriptions of relevant variables). Note that yardline is a real number in $\{0,1,...,99,100\}$ measuring the number of yards from the opponent’s endzone, where 00 represents a touchdown and $100$ indicates a safety. Holding out plays from 1999 to 2005 and from 2022 for various models and validation procedures discussed later in this paper, we are left with a primary dataset of $600,825$ plays, with $229,635$ first-down plays and $4,101$ non-tied games from 2006 to 2021, henceforth referred to as the “observed play-by-play football dataset.” We fit our win probability models using this dataset. The code for this study, which includes code to scrape the dataset used in this study, is publicly available on Github.222 https://github.com/snoopryan123/fourth_down
@@ 2.2 Example plays
The strength of a traditional decision recommendation is proportional to the effect size, or the estimated gain in win probability by making that decision. Notably, Baldwin, (2021) and Burke employ this decision procedure and post their fourth-down recommendations on X (formerly known as Twitter).333 Baldwin’s posts his fourth-down recommendations at @ben_bot_baldwin and Burke posts his at @bburkeESPN. Baldwin has posted a recommendation for many fourth-down plays since at least 2021444 As of the writing of this paper, @ben_bot_baldwin has over 11,000 posts (decision recommendations) on X. and Burke has posted a recommendation for select fourth down plays since at least 2022. We illustrate this procedure through two example plays.
Example play 1. First, we consider an example of Baldwin’s fourth-down decision charts. Figure 1 illustrates his charts for a fourth-down play in which the Patriots had the ball against the Colts in Week 10 of 2023. Baldwin views ${\mathsf{Go}}$ as a “strong” decision because he estimates that going for it provides a $3.6\%$ gain in win probability over attempting a field goal.
Example play 2. Next, we give an example of Burke’s decision chart. Figure 2 illustrates his chart for a play from the 2023 NFC Championship game. Burke views ${\mathsf{Go}}$ as the right decision because the yellow dot (denoting the actual play’s yardline and yards to go) lies squarely in the red region (denoting that ${\mathsf{Go}}$ is the estimated optimal decision) and is far from the decision boundary.
Figure 1: Baldwin’s decision charts for example play 1. The top image summarizes his models’ outputs for each decision, including estimated fourth-down win probability (in red), ${\mathsf{Go}}$/${\mathsf{FG}}$ success probability (second column), and win probability given ${\mathsf{Go}}$/${\mathsf{FG}}$ failure or success (third and fourth column). The bottom image is a screenshot of his post on X, which summarizes the recommended decision and the actual play.
Figure 2: Burke’s decision boundary chart for example play 2. The chart visualizes the estimated optimal decision according to effect size (color) as a function of yardline ($x$-axis) and yards to go ($y$-axis), holding the other game-state variables constant. The yellow dot denotes the actual play’s yardline and yards to go.
@@ 2.3 Estimating the win probability of a fourth-down decision
An offense has three possible decisions to make on fourth down: it can punt the ball, kick a field goal, or attempt a conversion (colloquially known as “going for it”), denoted ${\mathsf{Punt}}$, ${\mathsf{FG}}$, and ${\mathsf{Go}}$, respectively. In this section we estimate the win probability of a fourth-down decision in $\{{\mathsf{Punt}},{\mathsf{FG}},{\mathsf{Go}}\}$ in terms of first-down win probability (denoted $\mathsf{WP_{1}}$) and decision transition probabilities (punt outcome distribution, field goal success probability, and conversion outcome distribution). We estimate first-down win probability in Section 2.4 and decision transition probabilities in Section 2.5.
${\mathsf{Punt}}$ win probability. Suppose the offensive team has possession on fourth down at yardline $y$ and denote the remainder of the game-state by ${\mathbf{x}}$. If the offensive team punts, the opposing team has possession on first down at the subsequent yardline, which we model as a random variable. Hence the win probability of punting is one minus the opponent’s first-down win probability at the next yardline $y^{\prime}$ and next game-state ${\mathbf{x}}^{\prime}$,
We model the expected next yardline after punting as a function of yardline and punter quality in Section 2.5.
${\mathsf{FG}}$ win probability. We decompose field goal win probability in terms of field goal success probability and first-down win probability on the subsequent play,
[…]
We model the probability of a successful field goal as a function of yardline and kicker quality in Section 2.5. If the kicking team misses the field goal and the spot of the kick is within the 20 yardline, the opposing team has a first-down possession at yardline $80$. If the kicking team misses the field goal and the spot of the kick is beyond the 20 yardline, the opposing team has a first-down possession at the spot of the kick. The spot of the kick is typically 7 yards behind the line of scrimmage, so in terms of the current yardline $y$, the next yardline relative to the opposing team is $\min\{80,100-(y+7)\}$. Denoting the remaining game-state variables relative to the opposing team at the next play by ${\mathbf{x}}^{\prime}$, we have
[…]
If the kicking team makes the field goal, it scores three points and the opposing team has a first-down possession after a kickoff. Denoting the score differential relative to the kicking team by $s$, we have
${\mathsf{WP}}(\text{make }{\mathsf{FG}})=1-{\mathbb{E}}_{\text{kickoff}}[\mathsf{WP_{1}}(\text{yardline }y^{\prime},s^{\prime}=-s-3,{\mathbf{x}}^{\prime})].$ (6)
${\mathsf{WP}}(\text{make }{\mathsf{FG}})\approx 1-\mathsf{WP_{1}}(\text{yardline }{\mathbb{E}}_{\text{kickoff}}[y^{\prime}],\ s^{\prime}=-s-3,{\mathbf{x}}^{\prime}).$ (7)
${\mathsf{WP}}(\text{make }{\mathsf{FG}})\approx 1-\mathsf{WP_{1}}(\text{yardline }75,\ s^{\prime}=-s-3,{\mathbf{x}}^{\prime}).$ (8)
${\mathsf{Go}}$ win probability. Suppose the offensive team has possession on fourth down and $z$ yards-to-go at yardline $y$. If the offensive team goes for it and gains $\Delta\geq z$ yards, then in the next play it either has possession on first down at yardline $y-\Delta$ or scores a touchdown. Conversely, if the offensive team goes for it and gains $\Delta<z$ yards, then in the next play the opposing team has possession on first down at yardline $100-(y-\Delta)$. Hence the expected value of going for it on fourth down is
${\mathbb{E}}_{\text{go}}\big{[}\mathbb{I}\left(\Delta\geq z\right)\cdot\mathsf{WP_{1}}(\text{yardline }y-\Delta)+\mathbb{I}\left(\Delta<z\right)\cdot(1-\mathsf{WP_{1}}(\text{yardline }100-(y-\Delta)))\big{]}.$ (9)
This quantity is implicitly a function of the game-state.
This expectation is defined in terms of the conversion outcome conditional density $\allowbreak{\mathbb{P}}_{\text{go}}(\text{gain $\Delta$ yards}|{\mathbf{x}})$. This distribution is highly complex, non-normal, and non-symmetric. For instance, pass plays feature a spike at zero yards gained for incompletions. The mean yards gained and the tail vary as yardline changes. There is a spike at the endzone for touchdowns, which increases as the yardline approaches the endzone. The density differs for each team depending the particularities of a team’s play calling and personnel. Recent research has gone into estimating this conditional density: Biro and Walker, 2023a model the conditional density of a run play using a skew-$t$ distribution and Biro and Walker, 2023b model the conditional density of a pass play using a generalized gamma distribution, which they fit using Markov chain Monte Carlo methods.
For simplicity, we instead switch the order of the expectation and $\mathsf{WP_{1}}$ as we did in estimating ${\mathsf{Punt}}$ and ${\mathsf{FG}}$ win probability,
$\begin{split}&{\mathbb{P}}(\Delta\geq z|{\mathbf{x}})\cdot\mathsf{WP_{1}}(\text{yardline }y-{\mathbb{E}}_{\text{go}}[\Delta|{\mathbf{x}},\Delta\geq z],{\mathbf{x}}^{\prime})\\ +\ &{\mathbb{P}}(\Delta<z|{\mathbf{x}})\cdot(1-\mathsf{WP_{1}}(\text{yardline }100-(y-{\mathbb{E}}_{\text{go}}[\Delta|{\mathbf{x}},\Delta<z]),{\mathbf{x}}^{\prime})).\end{split}$ (10)
Here, ${\mathbf{x}}$ is the game-state on fourth down and ${\mathbf{x}}^{\prime}$ is the game-state on the subsequent play (which, as before, flips the game-state variables that are relative to the team with possession and doesn’t alter the other variables). We model the expected outcome of a successful conversion attempt ${\mathbb{E}}_{\text{go}}[\Delta|{\mathbf{x}},\Delta\geq z]$, that of a failed conversion attempt ${\mathbb{E}}_{\text{go}}[\Delta|{\mathbf{x}},\Delta<z]$, and conversion probability ${\mathbb{P}}(\Delta\geq z|{\mathbf{x}})$ in Section 2.5.
2.4 Estimating first-down win probability
@@
Now, we estimate first-down win probability as a function of game-state. Win probability estimates arise broadly from one of two classes of models, probabilistic state-space models or statistical models. On one hand, state-space models simplify the game of football into a series of transitions between game-states. Transition probabilities are estimated from play-level data and are then propagated into win probability by simulating games. When implemented correctly, these models are sensible ways to estimate ${\mathsf{WP}}$. However, they are difficult in practice, as they require: a careful encoding of the convoluted rules of football into a set of states and the actions between those states, careful estimation of transition probabilities, and enough computing power to run enough simulated games to achieve desired granularity. Each of these are nontrivial.
@@
On the other hand, statistical models are fit entirely from historical data. Given the results of a set of observed football plays, statistical models fit the relationship between certain game-state variables using data-driven regression or machine learning approaches. These models are widely used today in football analytics thanks to the accessbility of publicly available play-by-play data (e.g., nflFastR (Carl and Baldwin,, 2022)) and accessible off-the-shelf machine learning models (e.g., ${\mathsf{XGBoost}}$ (Chen and Guestrin,, 2016)). Additionally, due to a perceived abundance of data, flexible machine learning models are viewed as more “trustworthy” than previous mathematical models that make more simplifying assumptions. For these reasons, the open source win probability models used today in football analytics are statistical / machine learning models, which we focus on in this paper.
@@
The aforementioned models are fit from a historical play-by-play dataset that includes plays from all downs. Those models fit first-down win probability by including down as a covariate. We fit an ${\mathsf{XGBoost}}$ model just using first-down plays. This model predicts binary win/loss as a function of score differential, game seconds remaining, pre-game point spread, yardline, receive $2^{nd}$ half kickoff indicator, offensive team’s number of timeouts remaining, defensive team’s number of timeouts remaining, total score, and
scoreTimeRatio=score differential0.01+game seconds $\text{scoreTimeRatio}=\frac{\text{score differential}}{0.01+\text{game secondsremaining}}.$ (14)
Now, we compare the out-of-sample predictive performance of various first-down win probability models. Our full dataset consists of all football plays from 2006 to 2021. We split our dataset in half by randomly sampling 50% of the games. The first-down plays from the first 50% of these games form the hold-out test set. All the plays from the other 50% of these games form the training set. To tune the ${\mathsf{XGBoost}}$ models, we split the training set in half by randomly sampling 50% of the games from the training set. The plays from the first 50% of these games form the ${\mathsf{XGBoost}}$ training set, and the remaining plays form the validation set for hyper-parameter tuning. We then tune our ${\mathsf{XGBoost}}$ models in a similar fashion as Baldwin, (2021). We view the results of our prediction contest in Table 1. Our model fit from just first-down plays performs the best out-of-sample. Thus, in the remainder of this paper we use our proposed first-down win probability model.
model name model type plays in dataset out-of-sample ${\mathsf{logloss}}$
Our proposed model ${\mathsf{XGBoost}}$ first downs 0.440
Lock and Nettleton, (2014) ${\mathsf{RF}}$ all downs 0.447
Baldwin, (2021) ${\mathsf{XGBoost}}$ all downs 0.476
Table 1: Predictive performance of first-down win probability models.
2.5 Estimating fourth-down decision transition probabilities
Next, we estimate fourth-down decision transition probabilities. We model the expected outcome (yardline) of a punt, field goal success probability, conversion success probability, and the expected outcome (yardline) of a successful or a failed conversion attempt.
${\mathsf{Punt}}$ expected outcome model. We use linear regression to model the expected next yardline after a punt as a function of yardline and punter quality (denoted ${\mathsf{pq}}$, whose specification is detailed in Appendix B),
$\begin{split}{\mathbb{E}}_{\text{punt}}[\text{next yardline}]&=\ \vec{\alpha}\cdot\text{B-spline}\big{(}\text{yardline},\ \text{df}=4\big{)}+\ \beta_{1}\cdot{\mathsf{pq}}+\beta_{2}\cdot{\mathsf{pq}}\cdot\text{yardline}.\end{split}$ (15)
We use a B-spline to capture a nonlinear trend in yardline (de Boor,, 1978). The model is trained on a dataset of $36,493$ punts from 2006 to 2021, all beyond the 30 yardline. We visualize this model in Figure 3a. The above expectation is more properly defined in terms of the punt outcome conditional density $\allowbreak{\mathbb{P}}_{\text{punt}}(\text{next yardline}|{\mathbf{x}})$. But, that distribution is difficult to model (recall our discussion in Section 2.3 of modeling the conditional density of yards gained). A more sophisticated analysis would use that conditional density.
${\mathsf{FG}}$ success probability model. We use logistic regression to model the probability that a kicker makes a field goal as a function of yardline and kicker quality (denoted ${\mathsf{kq}}$, whose specification is detailed in Appendix B),
$\begin{split}\log\bigg{(}\frac{{\mathbb{P}}(\text{make }{\mathsf{FG}})}{1-{\mathbb{P}}(\text{make }{\mathsf{FG}})}\bigg{)}=\ \vec{\alpha}\cdot\text{B-spline}\big{(}\text{yardline},\ \text{df}=5\big{)}+\beta\cdot{\mathsf{kq}}.\end{split}$ (16)
Fitting this model on our dataset of $15,472$ observed field goals from 2006 to 2021 yields nontrivial probability predictions for extremely long field goals that have never before been made (e.g., nontrivial probability for a 73-yard field goal from the 55 yardline). To shrink these field goal probability predictions to zero, we impute 500 synthetic missed field goals, randomly distributed from the 51 to the 99 yardline, into our dataset. We visualize this model in Figure 3b.
Figure 3: (a) The expected next yardline after a punt ($y$-axis) according to our model as a function of yardline ($x$-axis) and punter quality (color). (b) The probability of making a field goal ($y$-axis) according to our model as a function of yardline ($x$-axis) and kicker quality (color).
${\mathsf{Go}}$ conversion probability model. Due to a small sample size of fourth-down conversion attempts, some existing fourth-down conversion probability models use third down as a proxy for fourth down (e.g., Romer, (2006)), as they are also high-pressure situations in which the offensive team attempts to obtain a first down. But, there may be a fundamental difference in conversion probability between third and fourth-down plays, perhaps due to psychological reasons or a differing distribution of play calls. In our model selection process, we train some models on a dataset consisting entirely of fourth-down plays and other models on a dataset consisting of third and fourth-down plays, testing models on a random 50% of fourth-down plays. The parameters of our best conversion probability model borrow strength from third-down plays.
We use logistic regression to model fourth-down conversion probability as a function of yards to go, down (third vs. fourth down), and our offensive and defensive quality metrics (whose specification is detailed in Appendix B): quarterback quality of the offensive team (${\mathsf{qbqot}}$), non-quarterback offensive quality of the offensive team (${\mathsf{oqrot}}$), defensive quality of the defensive team against the pass (${\mathsf{dqdtap}}$), and defensive quality of the defensive team against the run (${\mathsf{dqdtar}}$). We use these team quality metrics, rather than pre-game point spread, since they yield better predictive performance. This may be because quarterback quality significantly impacts conversion probability while the other team quality measures have a small impact, as seen in Figure 4. Formally, we model
$\begin{split}\log\bigg{(}\frac{{\mathbb{P}}({\text{convert}})}{1-{\mathbb{P}}({\text{convert}})}\bigg{)}=&\ \vec{\alpha_{1}}\cdot\mathbb{I}\left(\text{fourth down}\right)\cdot\text{B-spline}\big{(}\log(\text{yards to go}+1),\ \text{df}=4\big{)}\\ +&\ \vec{\alpha_{2}}\cdot\mathbb{I}\left(\text{third down}\right)\cdot\text{B-spline}\big{(}\log(\text{yards to go}+1),\ \text{df}=4\big{)}\\ +&\ \beta_{0}+\beta_{1}\cdot{\mathsf{qbqot}}+\beta_{2}\cdot{\mathsf{oqrot}}+\beta_{3}\cdot{\mathsf{dqdtap}}+\beta_{4}\cdot{\mathsf{dqdtar}}.\end{split}$ (17)
Figure 4: Fourth-down conversion probability ($y$-axis) according to our model as a function of yards to go ($x$-axis) and team quality (color). Quarterback quality has by far the largest impact on conversion probability, visualized by a larger spread amongst the colored lines.
${\mathsf{Go}}$ expected outcome models. We use linear regression to model the expected outcome (yards gained $\Delta$) of a conversion attempt with $z$ yards to go given that it is successful by
$\begin{split}{\mathbb{E}}_{\text{go}}[\Delta|{\mathbf{x}},\Delta\geq z]=&\ \vec{\alpha_{1}}\cdot\mathbb{I}\left(\text{fourth down}\right)\cdot\text{B-spline}\big{(}\log(\text{yards to go}),\ \text{df}=4\big{)}\\ +&\ \vec{\alpha_{2}}\cdot\mathbb{I}\left(\text{third down}\right)\cdot\text{B-spline}\big{(}\log(\text{yards to go}),\ \text{df}=4\big{)}\\ +&\ \vec{\alpha_{3}}\cdot\mathbb{I}\left(\text{yards to go}=1\right)\cdot\text{B-spline}\big{(}\text{yardline},\ \text{df}=3\big{)}\\ +&\ \vec{\alpha_{4}}\cdot\mathbb{I}\left(\text{yards to go}\neq 1\right)\cdot\text{B-spline}\big{(}\text{yardline},\ \text{df}=4\big{)}\\ +&\ \beta_{0}+\beta_{1}\cdot{\mathsf{qbqot}}+\beta_{2}\cdot{\mathsf{oqrot}}+\beta_{3}\cdot{\mathsf{dqdtap}}+\beta_{4}\cdot{\mathsf{dqdtar}}.\end{split}$ (18)
Similarly, we model the expected outcome of a conversion attempt with $z$ yards to go given that it is unsuccessful by
$\begin{split}{\mathbb{E}}_{\text{go}}[\Delta|{\mathbf{x}},\Delta<z]=&\ \alpha_{1}\cdot\mathbb{I}\left(\text{fourth down}\right)\cdot\log(\text{yards to go}+1)\\ +&\ \alpha_{2}\cdot\mathbb{I}\left(\text{third down}\right)\cdot\log(\text{yards to go}+1)\\ +&\ \beta_{0}+\beta_{1}\cdot{\mathsf{qbqot}}+\beta_{2}\cdot{\mathsf{oqrot}}+\beta_{3}\cdot{\mathsf{dqdtap}}+\beta_{4}\cdot{\mathsf{dqdtar}}.\end{split}$ (19)
These two expectations are more properly defined in terms of the conversion outcome conditional density $\allowbreak{\mathbb{P}}_{\text{go}}(\text{gain $\Delta$ yards}|{\mathbf{x}})$. But, as discussed in Section 2.3, that distribution is highly complex, non-normal, and non-symmetric, making it difficult to model. A more sophisticated analysis would use that conditional density.